OpenStudy (anonymous):

A ball is thrown upwards, from a building. It's height(h), in meters above the ground seconds(t) after being thrown is given by h(t)=-5t^2+15t+50 a) How tall is the building? (explain your answer) b) How long does it take the ball to reach the ground? c) For how long is the ball 55m above the ground?

OpenStudy (anonymous):

a) h(t)=-5t^2+15t+50 h(0)=0+0+50 h=50